OpenStudy (goformit100):

Which of the following is not an equivalent form of the compound inequality x + 15 less than or greater to 12 and x + 15 greater than or equal to 5 12 greater than or equal to x + 15 greater than or equal to 5 -3 greater than or equal to x greater than or equal to -10 A number line with a closed circle on -10, a closed circle on -3, and shading in between. A number line with a closed circle on -10, shading to the left, and a closed circle on -3, shading to the right.

OpenStudy (texaschic101):

does your inequality say " less then or greater " ??

OpenStudy (texaschic101):

Is that supposed to be less then or equal ?

OpenStudy (goformit100):

Which option is correct ?

OpenStudy (texaschic101):

less then or equal would be more correct then less then or greater

OpenStudy (texaschic101):

let me try it...hold on

OpenStudy (texaschic101):

x + 15 <= 12 x <= 12 - 15 x <= -3 x + 15 >= 5 x >= 5 - 15 x >= -10 the last one is incorrect I see you have changed your avatar.....it looks good :)